# Towards Principled Assessment of Tabular Data Synthesis Algorithms

### v1

@@ Appendix E Misleading conclusion
We find that some statements in data synthesis may be misleading or even incorrect due to the lack of principled evaluations. We highlight some of them as below:
@@
  • In [32], authors show that the machine learning performance on TabDDPM is even better than that on real data, which implies that synthetic data can be a perfect (even better) substitute for real data. However, this statement may be incorrect due to inadequate model tuning and improper data shuffling practices. Our evaluations show that even simple models, such as linear regression, can achieve better performance on real data than on high-quality synthetic data (i.e., generated by TabDDPM).
@@
  • Some studies [32, 34, 31, 30] prioritize machine learning efficacy as the primary (if not only) fidelity evaluation metric. This approach is problematic because data synthesis can be biased to label attributes, and a high machine learning efficacy score does not necessarily equate to high fidelity in synthetic data.
@@
  • Extensive studies [71, 34, 31, 32, 68] use Distance to Closest Records (DCR) to evaluate the privacy of synthetic data and assert their models are safe. However, in this paper we show that DCR fails to serve as an adequate measure of privacy. We also provide evidence that many recently introduced HP methods exhibit significant risks of membership disclosure, which are often ignored by HP synthesis algorithms.

### v2

@@ Appendix G Misleading conclusion
We find that some statements in data synthesis may be misleading or even incorrect due to limitations of evaluation metrics or methodologies. We highlight some of them as below:
@@
  • In (Kotelnikov et al., 2023), authors show that the machine learning performance on TabDDPM is even better than that on real data, which implies that synthetic data can be a perfect (even better) substitute for real data. However, this statement may be incorrect due to inadequate model tuning and improper data shuffling practices. Our evaluations show that even simple models, such as linear regression, can achieve better performance on real data than on high-quality synthetic data (i.e., generated by TabDDPM).
@@
  • Some studies (Kotelnikov et al., 2023; Lee et al., 2023; Kim et al., 2022; Jordon et al., 2018) prioritize machine learning efficacy as the primary (if not only) fidelity evaluation metric. This approach is problematic because data synthesis can be biased to label attributes, and a high machine learning efficacy score does not necessarily equate to high fidelity in synthetic data.
@@
  • Extensive studies (Zhao et al., 2021; Lee et al., 2023; Kim et al., 2022; Kotelnikov et al., 2023; Zhang et al., 2023b) use Distance to Closest Records (DCR) to evaluate the privacy of synthetic data and assert their models are safe. However, in this paper we show that DCR fails to serve as an adequate measure of privacy. We also provide evidence that many recently introduced HP methods exhibit significant risks of membership disclosure, which are often ignored by HP synthesis algorithms.
Appendix H Computing Wasserstein Distance
When the cost matrix is large and dense, computing this metric can be prohibitively time-consuming. Several options are provided to address this problem. (i) Sinkhorn distance (Cuturi, 2013) provides a fast approximation to the Wasserstein distance by penalizing the objective with an entropy term. (ii) Sliced-Wasserstein distance (Bonneel et al., 2015), which uses Radon transform to linearly project data into one dimension, can be efficiently computed. (iii) Reducing the size of the cost matrix by randomly sampling a small set of points from the probability densities. These computational strategies ensure the Wasserstein distance remains a feasible solution for fidelity evaluation, even for complex and high-dimensional marginals. In practice, when the domain of target probability distribution is large (i.e., $n\geq 5,000$), we find that randomly sampling half of the elements is both effective and efficient.